OpenStudy (anonymous):

Rewrite using the distributive property and then simplify (2x-3y)6

OpenStudy (matt101):

You first need to distribute the 6. That means you multiply each term in the brackets by 6: \[6(2x-3y)\]\[=6(2x)-6(3y)\]\[=12x-18y\]